The length of a rectangle is 4 less than twice the width. The area of the rectangle is 70 square feet. Find the width , w, of the rectangle algebraically
Answer by jorel555(1290)

l=2w-4;
w(2w-4)=70
2w^2-4w=70
w^2-2w=35
w^2-2w-35=0
(w-7)(w+5)=0
w=7,-5
The width of the rectangle is 7 feet!!!!!!!